Operations Report – February 27th

Crew 206 Operations Report 27 – FEB – 2019

SOL: 3

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need
replacement on generator. A contractor has been arranged to make these
repairs within a few days.

Generator (hours run): Turned off at 0800 (27th Feb), Turned on at 1840
(27th Feb)

Solar— SOC % – Turned off Gen (24th Feb) 64%, Turned on Gen (24th Feb) 59%

Diesel Reading – 3/4 (on a scale of 0 – 1/4 – 1/2 – 3/4 – 1)

Propane Reading – 21 (± 2)

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Water (Static tank) – around 500 gallons (90 %)

Water differential (Static tank) – 12 gallons (± 2 gallons)

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – 276.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01415960 units

Toilet tank emptied: Yes

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 52.4 hours

Beginning charge: 100%

Ending charge: 34%

Currently charging: Yes

Curiosity rover used: Used

Hours: 76.7 hours

Beginning charge: 100%

Ending charge: 61%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78.3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Opportunity’s battery decreased faster
under 60%, and lasted way less than Curiosity’s. (for more
detailed battery readings during the EVA, cf EVA report)

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Used for Staff

Staff Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Lost Internet access a lot again this
morning. Did a software reset on the modem. Working
without interruption now.

Summary of suits and radios: Suit #9 Fixed, currently charging. Suit #2
experienced fogging. Checked it, seems ok

Summary of Hab operations: Hotfixed the lamp bulb in upper
deck. It is stable now. Inspection revealed the plastic bulb is broken on
several loactions.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: Nothing to report

Summary of RAM operations:

Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing more to report.

Copyright © The Mars Society. All rights reserved. | Main Site